OpenStudy (anonymous):

A metal rod is 25.000cm long at 25.0 degrees Celsius. When heated to 102.0 degrees Celsius, it is 25.054 cm long. What is the coefficient of linear expansion for this metal?

OpenStudy (kropot72):

\[\alpha=\frac{1}{l}\times \frac{\Delta l}{\Delta T}\] where alpha is the coefficient of linear expansion, l is the original length and delta l is the change in length resulting from the change in temperature delta T. \[\alpha=\frac{1}{25.000}\times \frac{(25.054-25.000)}{(102-25)}=you\ can\ calculate\]

OpenStudy (anonymous):

Thank You ! :)

OpenStudy (kropot72):

You're welcome :)